# Artificial social influence via human-embodied AI agent interaction in immersive virtual reality (VR): Effects of similarity-matching during health conversations

### v1

@@ 5.1.1 Artificial Influence: A New Field of Research about How AI Can Change Human Behavior
Our research paradigm, then, augments existing research by illustrating how the persuasive power of factors such as similarity can vary based on the context (e.g., conversation about health vs. social dialogue; one-time vs. longitudinal interaction) and other variables. With this in mind, it is also clear that our analyses have barely scratched the surface of what is possible in terms of studying the actual interaction processes between humans and embodied AIs - in an over-time fashion and simultaneously monitoring effects on multiple social behaviors (e.g. gaze, facial expressions, verbal content, etc.).
@@ 6 Conclusion
To summarize, we examined the influence of gender-matched VR-embodied conversation agents (ECA). This was achieved by combining the potential of VR to simulate human-like personae (with a physical appearance, a voice, and lively nonverbal dynamics) with a state-of-the-art LLM-based AI agent. Regarding the impact of embodiment and VR, we find that participants experienced greater immersion and presence after interacting with VR-ECA vs. text-based CAs. The similarity manipulation showed that female participants liked the VR-ECA more than male participants, regardless of the agents’ gender. In addition, opposite-gender VR-ECAs enhanced healthy snack selection and gaze duration. Overall, we expect that the massive expansion of AI-LLMs is soon going to merge with VR, creating a fully embodied human-like simulacra. As this happens, current text-based large-language models (tb-LLMs) will evolve into multimodal communication models (MCMs) that feature the full spectrum of social embodiment. Although such a development will certainly have social consequences beyond our current research, from a research standpoint, this will enable us to examine a myriad of questions related to persuasion and social interaction - far beyond the basic gender-matching effects studied here.

### v2

@@ 5.1.1 Artificial Influence: A New Field of Research about How AI Can Change Human Behavior
Our research paradigm, then, augments existing research by illustrating how the persuasive power of factors such as interpersonal similarity can vary based on the context (e.g., conversation about health vs. social dialogue; one-time vs. longitudinal interaction) and other variables. With this in mind, it is also clear that our analyses have barely scratched the surface of what is possible in terms of studying the actual interaction processes between humans and embodied AIs - in an over-time fashion and simultaneously monitoring effects on multiple social behaviors (e.g. gaze, facial expressions, verbal content, etc.).
@@ 6 Conclusion
To summarize, we examined the influence of gender-matched VR-embodied conversation agents (ECA). This was achieved by combining the potential of VR to simulate human-like personae (with a physical appearance, a voice, and lively nonverbal dynamics) with a state-of-the-art LLM-based AI agent. Regarding the impact of embodiment and VR, we find that participants experienced greater immersion and presence after interacting with VR-ECA vs. text-based CAs. The similarity manipulation showed that female participants liked the VR-ECA more than male participants, regardless of the agents’ gender. In addition, opposite-gender VR-ECAs showed the potential to enhance overt attention and healthy snack selection. Overall, we expect that the massive expansion of AI-LLMs is soon going to merge with VR, creating a fully embodied human-like simulacra. As this happens, current text-based large-language models (tb-LLMs) will evolve into multimodal communication models (MCMs) that feature the full spectrum of social embodiment. Although such a development will certainly have social consequences beyond our current research, from a research standpoint, this will enable us to examine a myriad of questions related to persuasion and social interaction - far beyond the basic gender-matching effects studied here.